An airplane circle the earth at 0.6c. how many years must elapse before a clock in the plane and the one on the ground differ by 1second(0.694S)


1
Expert's answer
2021-07-13T10:42:47-0400

Let T denote the elapsed time on the ground, and t the elapsed time on the airplane. Then we have 


"T=\frac{t}{\sqrt{1-\frac{v^2}{c^2}}}\\T-t=1\\
T=\frac{T-1}{\sqrt{1-0.6^2}}\\T=5\ s"